OpenStudy (crashonce):

An AP consists of an even number of terms. The sums of the even and odd numbered terms are 300 and 240 respectively. If the first term is less than the last term by 118.5, the number of terms in the AP is?

OpenStudy (crashonce):

@ganeshie8 @jim_thompson5910

OpenStudy (crashonce):

@wio

OpenStudy (shinalcantara):

Sn = 300 + 240 Sn = 540 --------------- An = A1 + (n-1)d An - A1 = (n-1)d An-A1 = 118.5 118.5 = (n-1)d d = 118.5/(n-1) ------equation 1 --------------- Sn = (2A1 + (n-1)d)n/2 540(2) = (2A1 + (n-1)d)n 1080 = [2A1 + (n-1)(118.5)/(n-1)]n 1080 = [2A1 + 118.5]n n = 1080/[2A1 + 118.5] --------------- I was stuck with A1 and n as the remaining variables =[

OpenStudy (crashonce):

ok @wio ideas?